What is the x-intercept for the line 2x+1=y?
Answer by stanbon(48558)

x-intercept for the line 2x+1=y
-----------
Let y=0
Then 2x+1=0
x=-1/2
The x-intercept is (-1/2,0)
cheers,
Stan H.